# Capacitive detection of the topological magnetoelectric effect

### v1

@@ Abstract
The TME can also be probed indirectly using transport. In the absence of disorder the TME of thin films with small surface magnetizations is equivalent to separate half-quantized surface state Hall conductivities of opposite signs on top and bottom film surfaces. Assuming that the two surfaces conduct in parallel, only one surface will contribute to the Hall conductivity when only one surface is magnetized. Its Hall conductivity can then be extracted by inverting the measured total resistivity. Indeed, experimental observation of a near half-quantized Hall effect has been reported in a thin film with magnetization on one surface only [37, 38]. This measured surface Hall conductivity implies that charge flows into or out of that surface as magnetic field is varied. Similarly, measurements of the total QAH effect [39, 40] in thin films when the magnetization on the two surfaces are parallel can be viewed as providing information that is equivalent to that obtained by measurements in a device with only one magnetized surface. The Hall conductivity is simply twice as large as in the single-magnetic-surface case. The accuracy of quantization and the finite plateau width of the anomalous Hall effect, which is critical to identify quantized values, is assisted by disorder which makes states at the Fermi level in the 2-dimensional bulk irrelevant as long as they are localized. These surface localized states are not irrelevant for the TME in real disordered devices however, since they are related to the equilibrium charge transfer between layers after allowing the system to re-equilibrate following any change in the magnetic field value. The time scale required for this re-equilibration can be estimated by measuring the frequency-dependence of the capacitance in oscillating magnetic fields. The TME quantization should be more accurate when the magnetic field frequency exceeds the surface equilibration rate.
@@
Acknowledgements— We acknowledge fruitful discussions with Cui-Zu Chang, Lingjie Zhou, Deyi Zhuo, and Xiaoda Liu. This work was supported by Welch Foundation grant F-2112 and by the Simons Foundation. C.M.C. acknowledges support from the Swedish Research Council (VR 2021-046229).
Supplementary Material for “Capacitive Detection of Topological Magneto-electric Effect”
The Hamiltonian used in this work is constructed using Dirac cones that are coupled by hybridizing states:
$\hat{H}=\sum_{{\mathbf{k}}_{\perp},ij,\mu\gamma}\Big{[}\Big{(}\,(-1)^{i}\hbar v_{D}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}+m_{i}\sigma_{z}\Big{)}_{\mu,\gamma}\delta_{ij}+\Delta_{ij}(1-\delta_{ij})\delta_{\mu,\gamma}\Big{]}\hat{c}_{{\mathbf{k}}_{\perp},i,\mu}^{\dagger}\hat{c}_{{\mathbf{k}}_{\perp},j,\gamma}.$ (7)
Here we have chosen $\hat{\bm{z}}$ as the stacking axis of topological insulator quintuple layers, ${\mathbf{k}}_{\perp}=(k_{x},k_{y})$, $i,j\in\{0,\ldots,2N-1\}$ where $i=2(l-1)$ $(i=2l-1)$ labels states that are associated with the bottom (top) surface of layer $l\in\{1,\ldots,N\}$, $\mu,\gamma\in\{\uparrow,\downarrow\}$ are spin labels, $\Delta_{ij}$ are the intra- and inter-layer hybridization parameters, $v_{{}_{D}}$ is the velocity of an isolated Dirac cone, and $m_{i}$ are mass terms that account for exchange coupling to static local magnetic moments. The mass terms $m_{i}$ are related to the magnetizations via the equation $m_{i}=\sum_{l}J_{il}M_{l}$, where $M_{l}=0$ for non-magnetic layers and ±$\pm 1$ for magnetic layers; the Dirac-like states at the top or bottom surface of layer $l$ are coupled to the static magnetization of that layer (via $J_{\text{S}}$) and to that of the adjacent layer nearest that surface (via $J_{\text{D}}$), and hybridized only to adjacent surfaces, with hopping $\Delta_{\text{S}}$ for opposite surfaces of the same layer and $\Delta_{\text{D}}$ for the closest surface of adjacent layers. When set the parameter $J_{D}=J_{S}$, this model describe the non-magnetic topological insulators with a surface magnetization $J_{S}$.
In the presence of an external magnetic field ${\mathbf{B}}=-B\hat{\bm{z}}$, the $\bm{k}\cdot\bm{p}$ model is mapped to an envelope function approximated Hamiltonian that is minimally-coupled to the electromagnetic field: $\hbar\bm{k}\rightarrow-i\hbar\bm{\nabla}-(eBx/c)\hat{\bm{y}}$ in the Landau gauge. Here $e$ is the elementary charge unit. Applying this recipe to Eq. (7) and seeking energy eigenvectors of the form $\Psi_{E,q_{y}}(x,y)=e^{iq_{y}y}\Phi_{E,q_{y}}(x)/\sqrt{L_{y}}$, where $\Phi_{E,q_{y}}(x)$ has $2N\times 2$ components that account for layer and spin label degrees of freedom,
$H_{D}=\hbar v_{{}_{D}}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}\to\hbar\omega_{c}\begin{pmatrix}0&a^{\dagger}\\ a&0\end{pmatrix},$ (8)
where $a=\frac{1}{\sqrt{2}}(\tilde{x}+\partial_{\tilde{x}})$ and $a^{\dagger}=\frac{1}{\sqrt{2}}(\tilde{x}-\partial_{\tilde{x}})$ are Landau level raising and lowering (differential) operators, $\omega_{c}\equiv v_{D}/\sqrt{2}l_{B}$, $l_{B}\equiv\sqrt{\hbar c/eB}$ is the magnetic length, and $\tilde{x}\equiv l_{B}q_{y}-x/l_{B}$. The density of states in the presence of magnetic field is:
$D(E)=\frac{eB}{h}\sum_{l=0}^{\infty}{\delta(E-E_{l})},$ (9)
where $E_{l}$ is the Landau level energies. Considering the disorder effect, the density of states for Landau levels are broadened as:
$D(E)=\frac{eB}{\pi h}\sum_{l=0}^{\infty}{\frac{\Gamma}{\Gamma^{2}+(E-E_{l})^{2}}}.$ (10)
Here we considered a Lorentzian broadening where $\Gamma$ is a parameter specifying the width which is determined by the strength of disorder. The projected integrated density of states to the top and bottom surface defined in the main text are calculated by
[…]
in which $N$ is the thickness of thin films, $E_{F}$ is the Fermi level, and $i$ labels the Dirac cones locations. The $\delta N(E)\equiv N_{t}(E_{F})-N_{b}(E_{F})$ is then used the characterize the electric polarization.

### v2

@@ Abstract
The TME can also be probed indirectly using transport. In the absence of disorder the TME of thin films with small surface magnetizations is equivalent to separate half-quantized surface state Hall conductivities of opposite signs on top and bottom film surfaces. Assuming that the two surfaces conduct in parallel, only one surface will contribute to the Hall conductivity when only one surface is magnetized. Its Hall conductivity can then be extracted by inverting the measured total resistivity. Indeed, experimental observation of a near half-quantized Hall effect has been reported in a thin film with magnetization on one surface only [42, 43]. This measured surface Hall conductivity implies that charge flows into or out of that surface as magnetic field is varied. Similarly, measurements of the total QAH effect [44, 45] in thin films when the magnetization on the two surfaces are parallel can be viewed as providing information that is equivalent to that obtained by measurements in a device with only one magnetized surface. The Hall conductivity is simply twice as large as in the single-magnetic-surface case. The accuracy of quantization and the finite plateau width of the anomalous Hall effect, which is critical to identify quantized values, is assisted by disorder which makes states at the Fermi level in the two-dimensional bulk irrelevant as long as they are localized. These surface localized states are not irrelevant for the TME in real disordered devices, however, since they are related to the equilibrium charge transfer between layers after allowing the system to reequilibrate following any change in the magnetic field value. The time scale required for this reequilibration can be estimated by measuring the frequency dependence of the capacitance in oscillating magnetic fields. The TME quantization should be more accurate when the magnetic field frequency exceeds the surface equilibration rate. Ultimately, material platforms capable of hosting the TME that are presently available always have some degree of disorder. Measurements of the type we propose in this Letter may help provide some guidance to materials efforts that are designed to reduce the disorder in topological insulator surface states.
@@
Acknowledgments— We acknowledge fruitful discussions with Cui-Zu Chang, Lingjie Zhou, Deyi Zhuo, Xiaoda Liu and Shengwei Chi. This work was supported by the Robert A. Welch Foundation under Grant Welch F-2112 and by the Simons Foundation. C.M.C. acknowledges support from the Swedish Research Council (VR 2021-046229).